# SpiderCat: A Catalog of Compact Binary Millisecond Pulsars

### v1

@@ 3.6 Spin-orbit
The six BW outliers in Fig. 9 (panel b) are PSR J1745$+$1017, PSR J2234$+$0944, PSR J0610--2100, PSR J1630$+$3550, PSR J2214$+$3000, and PSR J0541$+$2959g. All exhibit low spin-down luminosities ($\dot{E}<2\times 10^{34}$ erg/s; no estimate exists for PSR J0541$+$2959g due to an unmeasured $\dot{P}$), consistent with Group 2 sources. Therefore, their higher orbital periods are unlikely to result from irradiation effects. Notably, three of these sources – PSRs J0610--2100, J2214$+$3000, and J2234$+$0944 – are the only spiders included in pulsar timing arrays (NANOGrav: Z. Arzoumanian et al. 2018; EPTA: G. Desvignes et al. 2016; A.-S. Bak Nielsen et al. 2020), demonstrating remarkably stable timing over a decade (E. van der Wateren et al., 2022). This is atypical, as most spiders show timing irregularities on timescales of months to a few years (e.g., Z. Arzoumanian et al., 1994; H. J. Pletsch & C. J. Clark, 2015; J. S. Deneva et al., 2016; G. Shaifullah et al., 2016; C. J. Clark et al., 2020), likely caused by changes in the companion star’s gravitational quadrupole moment (J. H. Applegate & J. Patterson, 1987; J. H. Applegate & J. Shaham, 1994) or tidal forces from asynchronous rotation (A. D. van Staden & J. Antoniadis, 2016). While not unique among BWs, these outliers show no or only intermittent, low-frequency radio eclipses (E. D. Barr et al., 2013b; E. F. Lewis et al., 2023; P. F. Wang et al., 2025; S. Kumari et al., 2025), suggesting either an absence or minimal presence of ionized material in the system, or that such material does not intersect the line of sight. One possibility is that the companions are semi-degenerate or degenerate stars, which are more resistant to ablate. This interpretation aligns with the idea that, like the RB outliers, these BW outliers host atypical companions.
@@ 4 Conclusions
The study of compact binary MSPs is experiencing a remarkable surge, driven by an increasing number of discoveries and their importance for fundamental astrophysics. This rapidly growing class of pulsars, known as “spiders”, consists of neutron stars in tight binary orbits that gradually strip material from their low-mass companions. These systems are among the best laboratories for studying pulsar evolution and are key to identifying the most massive neutron stars, which provide crucial constraints on the equation of state of ultra-dense matter.
@@
The number of known spider systems and candidates in the Galactic field has now surpassed 100, marking the transition into an era of population-level statistical studies. In addition to the two main subclasses, several peculiar systems and subclasses have emerged, including mildly recycled pulsars with slow spin periods, transitional systems that alternate between accretion-powered and rotation-powered states (tMSPs), and those with unusual companion stars, such as extremely low-mass companions (“tidarren”) or giant companions (“huntsman”).
@@
In this paper, we compiled and summarized multiwavelength data and radio timing parameters for the current population of spider systems. The paper is accompanied by a public online database (https://astro.phys.ntnu.no/SpiderCAT), which we aim to update continuously as new sources are discovered or existing entries are refined [^12] We welcome community feedback, see https://astro.phys.ntnu.no/SpiderCAT/acknowledgements.. The catalog highlights the most relevant parameters of each system, as presented here.
Utilizing the catalog, we found that:
@@
  • Confirmed spiders (excluding mildly recycled systems) have spin periods between 1.4 and 5.4 ms and spin period derivatives ranging from $10^{-21}$ to $7.5\times 10^{-20}$ s/s.
@@
  • Excluding tidarren and huntsman systems, orbital periods range from approximately 2 to 26 hours. Only four RBs (12.5%) have orbital periods shorter than 4 hours, compared to 27 BWs (55%).
@@
  • Parameters observed exclusively in RB systems include minimum companion masses greater than $M_{\rm c,min}>0.06,M_{\odot}$, X-ray luminosities exceeding $L_{X}>1.3\times 10^{32}$ erg/s, and $\gamma$-ray to X-ray flux ratios below $\log(F_{\gamma}/F_{X})<2.3$.
@@
  • In a color-magnitude diagram, most spiders lie between the main sequence and the white dwarf branch. In the Pan-STARRS filters, this region is approximately bounded from above by $M_{g}=4.5\times(g-r)+4.5$ and from below by $M_{g}=2.7\times(g-r)+10.3$, providing a useful criterion for identifying spider candidates in optical surveys. Notably, RBs exhibit absolute $g$-band magnitudes roughly two magnitudes brighter than BWs at similar colors.
@@
  • The spider distribution is peaked toward the Galactic plane (Fig. 2). We estimate a Galactic scale height scale (Fig. 3) of $z_{e}=0.73\pm 0.15$ kpc for spiders.
@@
The future of spider searches remains bright. Many spiders are likely hidden in the Galactic plane – a region historically underexplored due to source confusion (with the notable recent exception of the FAST-GPPS survey). Given the relatively high Galactic scale height of spiders, many may also remain undetected at higher latitudes. Additionally, a large number of unassociated Fermi sources remain to be classified ($>$ 2000; S. Abdollahi et al., 2022a; M. G. F. Mayer & W. Becker, 2024), many of which may host a spider pulsar. Therefore, we anticipate the number of known Galactic field spiders to continue increasing.
@@
We thank Iacob Nedreaas and Bogdan Voaidas for software contributions to the catalog’s search engine and web server. We also thank Devina Misra for discussion on the binary evolutionary models of spiders. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352, PI: M. Linares). This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The national facility capability for SkyMapper has been funded through ARC LIEF grant LE130100104 from the Australian Research Council, awarded to the University of Sydney, the Australian National University, Swinburne University of Technology, the University of Queensland, the University of Western Australia, the University of Melbourne, Curtin University of Technology, Monash University and the Australian Astronomical Observatory. SkyMapper is owned and operated by The Australian National University’s Research School of Astronomy and Astrophysics. The survey data were processed and provided by the SkyMapper Team at ANU. The SkyMapper node of the All-Sky Virtual Observatory (ASVO) is hosted at the National Computational Infrastructure (NCI). Development and support of the SkyMapper node of the ASVO has been funded in part by Astronomy Australia Limited (AAL) and the Australian Government through the Commonwealth’s Education Investment Fund (EIF) and National Collaborative Research Infrastructure Strategy (NCRIS), particularly the National eResearch Collaboration Tools and Resources (NeCTAR) and the Australian National Data Service Projects (ANDS). This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. This research has made use of data obtained from the 4XMM XMM-Newton serendipitous source catalogue compiled by the XMM-Newton Survey Science Centre consortium. {contribution} KK led the development of this work, including writing and preparing the original manuscript draft, designing and implementing the software tools, curating and analyzing the dataset, generating the visualizations, and conducting the formal analysis. KK was also responsible for the submission process and the overall coordination of manuscript preparation. ML conceived the original idea for the project and contributed to the spider definition, data curation and software development. ML reviewed, edited and contributed to the manuscript throughout its preparation, secured the funding necessary to support the research and manages the project’s git repository.

### v4

@@ 3.6 Spin-orbit
A few BWs can be found at intermediate spins (3–4 ms) and orbital periods (6–10 hr) offset from most sources. These are PSR J2234++0944, PSR J0610-2100, PSR J1630++3550, PSR J2214++3000, and PSR J0541++2959g. All exhibit low spin-down luminosities ($\dot{E}<2\times 10^{34}$ erg s-1; no estimate exists for PSR J0541++2959g due to an unmeasured $\dot{P}$), in line with Cluster 2 sources. Therefore, their higher orbital periods are unlikely to result from irradiation effects. Notably, three of these sources—PSRs J0610-2100, J2214++3000, and J2234++0944—are the only spiders included in pulsar timing arrays (NANOGrav: Z. Arzoumanian et al. 2018; EPTA: G. Desvignes et al. 2016; A.-S. Bak Nielsen et al. 2020), demonstrating remarkably stable timing over a decade (E. van der Wateren et al., 2022). This is atypical, as most spiders show timing irregularities on timescales of months to a few years (e.g., Z. Arzoumanian et al., 1994; H. J. Pletsch & C. J. Clark, 2015; J. S. Deneva et al., 2016; G. Shaifullah et al., 2016; C. J. Clark et al., 2021), likely caused by changes in the companion star’s gravitational quadrupole moment (J. H. Applegate & J. Patterson, 1987; J. H. Applegate & J. Shaham, 1994) or tidal forces from asynchronous rotation (A. D. van Staden & J. Antoniadis, 2016). While not unique among BWs, these outliers show no or only intermittent, low-frequency radio eclipses (E. D. Barr et al., 2013; E. F. Lewis et al., 2023; S. Kumari et al., 2025; P. F. Wang et al., 2025), suggesting either an absence or minimal presence of ionized material in the system, or that such material does not intersect the line of sight. One possibility is that the companions are semidegenerate or degenerate stars, which are more resistant to ablate.
@@ 4 Conclusions
The study of compact binary MSPs is experiencing a remarkable surge, driven by an increasing number of discoveries and their importance for fundamental astrophysics. This rapidly growing class of pulsars, known as spiders, consists of neutron stars in tight binary orbits that gradually strip material from their low-mass companions. These systems are among the best laboratories for studying pulsar evolution and are key to identifying the most massive neutron stars, which provide crucial constraints on the equation of state of ultradense matter.
@@
The number of known spider systems and candidates in the Galactic field has now surpassed 100, marking the transition into an era of population-level statistical studies. In addition to the two main subclasses, several peculiar systems and subclasses have emerged, including mildly recycled pulsars with slow spin periods, transitional systems that alternate between accretion-powered and rotation-powered states (tMSPs), and those with unusual companion stars, such as extremely low-mass companions (tidarren) or giant companions (HM).
@@
In this paper, we compiled and summarized multiwavelength data and radio timing parameters for the current population of spider systems. The paper is accompanied by a public online database,  [^14] https://astro.phys.ntnu.no/SpiderCAT which we aim to update continuously as new sources are discovered or existing entries are refined. [^15] We welcome community feedback, see https://astro.phys.ntnu.no/SpiderCAT/acknowledgements. Tables 2–5 are also available in machine-readable form in Vizier. [^16] https://cdsarc.cds.unistra.fr/viz-bin/cat/J/ApJ/994/8 The catalog highlights the most relevant parameters of each system, as presented here.
Utilizing the catalog, we found the following.
  1. 1.
@@
    Confirmed spiders (excluding mildly recycled systems) have spin periods between 1.4 and 5.4 ms and spin period derivatives ranging from $10^{-21}$ to $7.5\times 10^{-20}$ s s-1.
  2. 2.
@@
    Excluding tidarren and HM systems, the orbital periods range from approximately 2 to 26 hr. Only four RBs (12.5%) have orbital periods shorter than 4 hr, compared to 27 BWs (55%).
  3. 3.
@@
    Parameters observed exclusively in RB systems include minimum companion masses greater than $M_{\rm c,min}>0.06,M_{\odot}$, X-ray luminosities exceeding $L_{X}>1.3\times 10^{32}$ erg s-1, and $\gamma$-ray to X-ray flux ratios below $\log(F_{\gamma}/F_{X})<2$.
  4. 4.
@@
    In a color-magnitude diagram, most spiders lie between the main sequence and the WD branch. In the Pan-STARRS filters, this region is approximately bounded from above by $M_{g}=4.5\times(g-r)+4.5$ and from below by $M_{g}=2.7\times(g-r)+10.3$, providing a useful criterion for identifying spider candidates in optical surveys. Notably, RBs exhibit absolute $g$-band magnitudes roughly 2 mag brighter than BWs at similar colors.
  5. 5.
@@
    The spider distribution is peaked toward the Galactic plane (Figure 2). We estimate a Galactic scale height scale (Figure 3) of $z_{e}=0.73\pm 0.15$ kpc for spiders.
@@
The future of spider searches remains bright. Many spiders are likely hidden in the Galactic plane—a region historically underexplored due to source confusion (with the notable recent exception of the FAST-GPPS survey). Given the relatively high Galactic scale height of spiders, many may also remain undetected at higher latitudes. Additionally, a large number of unassociated Fermi sources remain to be classified (>> 2000; S. Abdollahi et al., 2022; M. G. F. Mayer & W. Becker, 2024), many of which may host a spider pulsar. Therefore, we anticipate the number of known Galactic-field spiders to continue increasing.
@@
We thank Iacob Nedreaas and Bogdan Voaidas for software contributions to the catalog’s search engine and web server. We are also grateful to Devina Misra for discussions on the binary evolutionary models of spiders, and to the anonymous referee for comments that improved the paper. In addition, we thank Clément Vidal for pointing out an error in the Gaia DR3 ID numbers. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352, PI: M. Linares). This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The national facility capability for SkyMapper has been funded through ARC LIEF grant LE130100104 from the Australian Research Council, awarded to the University of Sydney, the Australian National University, Swinburne University of Technology, the University of Queensland, the University of Western Australia, the University of Melbourne, Curtin University of Technology, Monash University and the Australian Astronomical Observatory. SkyMapper is owned and operated by The Australian National University’s Research School of Astronomy and Astrophysics. The survey data were processed and provided by the SkyMapper Team at ANU. The SkyMapper node of the All-Sky Virtual Observatory (ASVO) is hosted at the National Computational Infrastructure (NCI). Development and support of the SkyMapper node of the ASVO has been funded in part by Astronomy Australia Limited (AAL) and the Australian Government through the Commonwealth’s Education Investment Fund (EIF) and National Collaborative Research Infrastructure Strategy (NCRIS), particularly the National eResearch Collaboration Tools and Resources (NeCTAR) and the Australian National Data Service Projects (ANDS). This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. This research has made use of data obtained from the 4XMM XMM-Newton serendipitous source catalogue compiled by the XMM-Newton Survey Science Centre consortium. {contribution}$K.$ led the development of this work, including writing and preparing the original manuscript draft, designing and implementing the software tools, curating and analyzing the data set, generating the visualizations, and conducting the formal analysis. $K.$ was also responsible for the submission process and the overall coordination of manuscript preparation. M.L. conceived the original idea for the project and contributed to the spider definition, data curation and software development. M.L. reviewed, edited, and contributed to the manuscript throughout its preparation, secured the funding necessary to support the research and manages the project’s GitHub repository.